# MLLM Machine Unlearning via Visual Knowledge Distillation

### v1

@@ 4.1.1 Datasets and Metrics
4.1.2 Models and Baselines
MLLMs: We conduct experiments using open-source MLLMs as base models. Specifically, we use LLaVA-1.5-7B-hf and Qwen2-VL-7B-Instruct, which serve as the foundation for our unlearning experiments. The vanilla models are trained following the official implementations provided by MLLMU-Bench and CLEAR. For each vanilla model, we perform selective unlearning of visual knowledge according to our proposed framework, while freezing the LLM weights.
Baselines: We compare our method with six baseline approaches: GA [thudi2022unrolling], which applies gradient ascent on the forget VQA set $D_{f}$; GA_Diff [liu2022continual], an improved GA variant with a joint loss to balance the forget set $D_{f}$ and retain set $D_{r}$; KL_Min [maini2024tofu], which aligns predictions on $D_{r}$ with the original model while maximizing KL divergence with $D_{f}$; NPO [zhang2024negative], which treats $D_{f}$ as negative-preference data and integrates unlearning into a preference optimization framework with an oracle model fine-tuned on $D_{r}$; MANU [liu2025modality], which introduces modality-aware neuron unlearning by selectively pruning neurons that contribute most to the forget set, preserving model performance while achieving balanced forgetting across modalities; MMUnlearner [huo2025mmunlearner], which uses saliency-weighted updates to selectively protect non-target parameters while forgetting visual knowledge.
All implementations are based on official MLLMU-Bench and CLEAR code, using the same learning rate, forgetting rounds, batch size, and full-precision training on four NVIDIA 4090 GPUs.
@@ 4.1.2 Models and Baselines
4.2 Main Results and Analysis
4.2.1 Effectiveness of Unlearning
On MLLMU-Bench, our method achieves the lowest accuracy on the Forget VQA task for both LLaVA-7B and Qwen2-VL, surpassing the state-of-the-art method MMUnlearner by 1.4% and 1.2%, respectively. Similarly, on CLEAR, our approach consistently surpasses MMUnlearner by 3.5% and 1.3%, respectively. This illustrates that our approach can effectively erase the visual knowledge associated with the target entity.
Our method also achieves state-of-the-art performance on both the Retain QA task and the Forget QA task, particularly in MLLMU-Bench dataset. This demonstrates that our approach effectively preserves the textual knowledge associated with both target and non-target entities.
Compared to textual knowledge forgetting, retaining the visual knowledge associated with non-target entities remains the most challenging issue. From the results on the Retain VQA and Real-world VQA, our approach shows substantial improvement, outperforming both MANU and MMUnlearner. This demonstrates that our method possesses a strong advantage in balancing unlearning effectiveness and model utility with respect to visual knowledge.
We present several examples where our approach effectively removes target visual knowledge while preserving non-target knowledge, as shown in Appendix. We compare our approach with GA, NPO, and MANU. Our method successfully erases the target visual content. In contrast, GA often leads to partial forgetting and distorted textual responses, whereas NPO tends to over-forget, degrading both visual and textual modalities.
4.2.2 Effect of VKD
The Visual Knowledge Distillation (VKD) module is designed to preserve the visual knowledge associated with non-target entities. Its strength is controlled by the parameter $\beta$ in Eq. 3. As shown in Fig. 2, when $\beta=0$—that is, when the VKD term in Eq. 3 is omitted—the visual knowledge related to non-target entities is not effectively preserved, as evidenced by the notably low accuracy on the Retain VQA task. In contrast, as $\beta$ increases, the accuracy on the Retain VQA set rises sharply, indicating a substantial improvement in preserving the visual knowledge associated with non-target entities.
Interestingly, our VKD module also contributes to forgetting the visual knowledge associated with the target entity, even though it was not originally designed for this purpose. As shown in Fig. 2, as $\beta$ increases, the accuracy on the Forget VQA set also exhibits a noticeable decline. These results illustrate that our approach effectively disentangles the forgetting of target visual knowledge from the preservation of non-target visual knowledge.
Notably, the strength of VKD should not be set excessively high. We observe that as $\beta$ increases, the accuracy on the Real-world VQA set initially rises but subsequently declines, indicating that an overly strong VKD may harm the model’s overall utility. Finally, we set the optimal strength to $\beta=0.3$, which achieves a balanced trade-off between the output-level objective and the VKD objective.
Moreover, based on the accuracy results on the Forget, Retain, and Real-world QA sets, we observe that they are only slightly influenced by the VKD module. This demonstrates that our VKD design does not compromise the preservation of textual knowledge.
4.2.3 Balance between Forgetting and Preservation
@@
The parameter $\alpha$ controls the strength of target visual knowledge forgetting relative to the preservation of other forms of knowledge, including the target textual knowledge (corresponding to $\mathcal{L}^{f}_{qa}$) and both the visual and textual knowledge associated with non-target entities (corresponding to $\mathcal{L}^{r}_{vqa}$ and $\mathcal{L}^{r}_{qa}$). As shown in Fig. 3, as $\alpha$ increases, the accuracy on the Forget VQA set drops rapidly, indicating that the forgetting of target visual knowledge is significantly enhanced. Meanwhile, the accuracies on the Retain VQA and Real-world VQA sets initially rise but subsequently decline, reflecting the inherent conflict between knowledge forgetting and preservation. Nevertheless, a proper forgetting strength of $\alpha=1.25$ achieves a desirable balance between the two objectives.
As shown in Fig. 3, the accuracies on the Forget, Retain, and Real-world QA sets remain largely insensitive to variations in $\alpha$, indicating that our approach effectively disentangles textual knowledge from visual knowledge.
@@
Table 2: Comparison of Unlearning Efficiency. Average running time per epoch (in minutes) on an NVIDIA RTX 4090 GPU. The GPU memory consumption is compared across these methods.
4.2.4 Efficiency of Unlearning
By freezing the large language model (LLM) and performing lightweight fine-tuning on the visual module, our method exhibits significant efficiency advantages. We measure the average per-epoch unlearning time under different configurations and compare it with other baseline methods on an NVIDIA 4090 GPU. As shown in Table. 2, our design achieves an excellent balance between forgetting precision and computational efficiency. Freezing the LLM reduces the overall unlearning time by more than 50%, while maintaining strong unlearning effectiveness. The introduction of VKD adds only minor overhead.
4.3 Ablation Study
As shown in Table. 3, the baseline method (i.e., the “B” row) adopts a full fine-tuning strategy, wherein both the visual module and the LLM backbone are fine-tuned. When partial fine-tuning is excluded and only the VKD and masking schemes are applied (i.e., the “B+Mask+VKD” row), we observe that this configuration is inferior to our approach (i.e., the “F+Mask+VKD” row) in preserving textual knowledge, as evidenced by the lower accuracy on the Forget QA task. Moreover, its running time is significantly longer than that of our approach.
@@ 4.3 Ablation Study
Table 3: Ablation study on the Forget/Retain/Realworld sets. Lower is better ($\downarrow$) for Forget VQA accuracy; higher is better ($\uparrow$) for the others. “B” denotes the baseline, which fine-tunes the entire MLLM model, whereas “F” denotes freezing the LLM while fine-tuning only the visual module. “F+mask+VKD” represents our proposed approach.
@@ 4.4 Robustness of Unlearning
When conducting a relearning attack using a small portion of the forgetting data, we evaluate the extent to which the forgotten visual knowledge can be recovered after the attack. Specifically, we define a new metric, Accuracy Gap (AG), defined as the difference between the post-relearning accuracy and the accuracy of the original model, where the post-relearning accuracy refers to the final accuracy once relearning procedure has stably converged. The larger the AG, the more robust the unlearning algorithm is. In practice, we measure the AG using the accuracy on the Forget VQA task, i.e., by comparing the difference in Forget VQA accuracy before and after the relearning process.
We conduct relearning attacks using 10%, 20%, and 30% of the forgotten visual data. As shown in Table. 4, as the proportion of fine-tuning data used for the relearning attack increases, more forgotten knowledge is gradually recovered. However, with 20% of the relearning data, the Forget VQA accuracy shows only a slight recovery, increasing from 29.3% to 30.6% (i.e., AG=1.3%), demonstrating that our approach effectively forgets the undesired knowledge and remains robust against relearning attacks. More results are provided in Appendix.
@@ 5 Conclusion
In this paper, we propose an MLLM-oriented machine unlearning approach that disentangles visual and textual knowledge, handling them separately—specifically, erasing target visual knowledge while preserving textual knowledge. In addition to leveraging output-level supervision, we introduce the Visual Knowledge Distillation (VKD) scheme, which incorporates intermediate visual features from a teacher model as an auxiliary supervision signal. Interestingly, VKD not only facilitates forgetting of target visual knowledge but also aids in preserving visual knowledge of non-target entities. Moreover, we evaluate the robustness of our MLLM unlearning approach against relearning attacks, demonstrating that the forgotten visual knowledge cannot be easily recovered.

### v2

@@ 4.1 Datasets and Metrics
4.2 Models and Baselines
MLLMs: We conduct experiments using open-source MLLMs as base models. Specifically, we use LLaVA-1.5-7B-hf and Qwen2-VL-7B-Instruct, which serve as the foundation for our unlearning experiments. The vanilla models are trained following the official implementations provided by MLLMU-Bench and CLEAR.
Baselines: We compare our method with six baseline approaches: GA (Thudi et al., 2022), which applies gradient ascent on the forget VQA set $D_{f}$; GA_Diff (Liu et al., 2022), an improved GA variant with a joint loss to balance the forget set $D_{f}$ and retain set $D_{r}$; KL_Min (Maini et al., 2024), which aligns predictions on $D_{r}$ with the original model while maximizing KL divergence with $D_{f}$; NPO (Zhang et al., 2024), which treats $D_{f}$ as negative-preference data and integrates unlearning into a preference optimization framework with an oracle model fine-tuned on $D_{r}$; MANU (Liu et al., 2025), which introduces modality-aware neuron unlearning by selectively pruning neurons that contribute most to the forget set, preserving model performance while achieving balanced forgetting across modalities; MMU(nlearner) (Huo et al., 2025), which uses saliency-weighted updates to selectively protect non-target parameters while forgetting visual knowledge.
@@ 4.2 Models and Baselines
4.3 Main Results and Analysis
Effectiveness of Unlearning.
As shown in Table 1, our approach can significantly reduce the accuracy of Forget VQA task on MLLMU-Bench for both LLaVA-7B and Qwen2-VL, surpassing the state-of-the-art method MMUnlearner by 1.4% and 1.2%, respectively. Similarly, on CLEAR, our approach consistently surpasses MMUnlearner by 2.0% and 1.3%, respectively. This illustrates that our approach can effectively erase the visual knowledge associated with the target entity.
Our method also retains the accuracy of both the Retain QA task and the Forget QA task, particularly in MLLMU-Bench dataset. This demonstrates that our approach effectively preserves the textual knowledge associated with both target and non-target entities.
Compared to target visual knowledge forgetting, retaining the visual knowledge associated with non-target entities remains the most challenging issue. From the results on the Retain VQA and Real-world VQA, our approach shows substantial improvement, outperforming both MANU and MMUnlearner. This demonstrates that our method possesses a strong advantage in balancing unlearning effectiveness and model utility with respect to visual knowledge.
We present several examples for comparing our approach with GA, NPO, and MANU, as shown in Appendix. We find that our method can successfully erases the target visual content. In contrast, GA often leads to partial forgetting and distorted textual responses, whereas NPO tends to over-forget, degrading both visual and textual modalities.
@@ 4.3 Main Results and Analysis
Effect of VKD.
The VKD module is designed to preserve the visual knowledge associated with non-target entities. Its strength is controlled by the parameter $\beta$ in Eq. 3. As shown in Fig. 2, when $\beta=0$—that is, when the VKD term in Eq. 3 is omitted—the visual knowledge related to non-target entities is not effectively preserved, as evidenced by the notably low accuracy on the Retain VQA task. In contrast, as $\beta$ increases, the accuracy of the Retain VQA rises sharply, indicating a substantial improvement in preserving the visual knowledge associated with non-target entities.
Interestingly, our VKD module also contributes to forgetting the visual knowledge associated with the target entity, even though it was not originally designed for this purpose. As shown in Fig. 2, as $\beta$ increases, the accuracy of the Forget VQA also exhibits a noticeable decline. These results illustrate that our approach effectively disentangles the forgetting of target visual knowledge from the preservation of non-target visual knowledge.
Notably, the strength of VKD should not be set excessively high. We observe that as $\beta$ increases, the accuracy of the Real-world VQA initially rises but subsequently declines, indicating that an overly strong VKD may harm the model’s overall utility. Finally, we set the optimal strength to $\beta=0.3$, which achieves a balanced trade-off between the output-level objective and the VKD objective.
Moreover, based on the accuracy of the Forget QA, Retain QA, and Real-world QA, we observe that they are only slightly influenced by the VKD module. This demonstrates that our VKD design does not compromise the preservation of textual knowledge.
Balance between Forgetting and Preservation.
@@
Table 2: Comparison of Unlearning Efficiency. Average running time per epoch (in minutes) on NVIDIA RTX 4090 GPU. The GPU memory consumption is compared across these methods.
The parameter $\alpha$ controls the strength of target visual knowledge forgetting relative to the preservation of other forms of knowledge, including the target textual knowledge (corresponding to $\mathcal{L}^{f}_{qa}$) and both the visual and textual knowledge associated with non-target entities (corresponding to $\mathcal{L}^{r}_{vqa}$ and $\mathcal{L}^{r}_{qa}$). As shown in Fig. 3, as $\alpha$ increases, the accuracy of the Forget VQA drops rapidly, indicating that the forgetting of target visual knowledge is significantly enhanced. Meanwhile, the accuracies of the Retain VQA and Real-world VQA initially rise but subsequently decline, reflecting the inherent conflict between knowledge forgetting and preservation. Nevertheless, a proper forgetting strength of $\alpha=1.25$ achieves a desirable balance between the two objectives.
As shown in Fig. 3, the accuracies of the Forget QA, Retain QA, and Real-world QA remain largely insensitive to variations in $\alpha$, indicating that our approach effectively disentangles textual knowledge from visual knowledge.
Efficiency of Unlearning.
By freezing the LLM backbone and applying lightweight fine-tuning to the visual module, our method achieves notable efficiency gains. We measure the average per-epoch unlearning time under different configurations and compare with other baseline methods on NVIDIA 4090 GPU. As shown in Table 2, our approach achieves a favorable trade-off between unlearning efficacy and computational efficiency. Freezing the LLM reduces the overall unlearning time by more than 50%, while maintaining strong unlearning effectiveness. The introduction of VKD adds only minor overhead.
@@
Table 3: Ablation study. Lower is better ($\downarrow$) for Forget VQA accuracy; higher is better ($\uparrow$) for the others. “B” denotes the baseline, which fine-tunes the entire MLLM model, whereas “F” denotes freezing the LLM while fine-tuning only the visual module. “F+mask+VKD” represents our proposed approach.
@@
4.4 Ablation Study
As shown in Table 3, the baseline method (i.e., the “B” row) adopts a full fine-tuning strategy, wherein both the visual module and the LLM backbone are fine-tuned. When partial fine-tuning is excluded and only the VKD and masking schemes are applied (i.e., the “B+Mask+VKD” row), we observe that this configuration is inferior to our approach (i.e., the “F+Mask+VKD” row) in preserving textual knowledge, as evidenced by the lower accuracy on the Forget QA task. Moreover, its running time is significantly longer than that of our approach.
@@ 4.5 Robustness of Unlearning
When conducting a relearning attack using a small portion of the forgetting data, we evaluate the extent to which the forgotten visual knowledge can be recovered after the attack. Specifically, we define a new metric, Accuracy Gap (AG), defined as the difference between the post-relearning accuracy and the accuracy of the original model, where the post-relearning accuracy refers to the final accuracy once relearning procedure has stably converged. The smaller the AG, the more robust the unlearning algorithm is. In practice, we measure the AG using the accuracy on the Forget VQA task, i.e., by comparing the difference in Forget VQA accuracy before and after the relearning process.
We conduct relearning attacks using 10%, 20%, and 30% of the forgotten visual data. As shown in Table 4, as the proportion of fine-tuning data used for the relearning attack increases, more forgotten knowledge is gradually recovered. However, with 20% of the relearning data, the Forget VQA accuracy shows only a slight recovery, increasing from 29.3% to 30.6% (i.e., AG=1.3%), demonstrating that our approach effectively forgets the undesired knowledge and remains robust against relearning attacks. More results are provided in Appendix C.
@@ 5 Conclusion
In this paper, we propose an MLLM unlearning approach that fine-tunes only the visual module while keeping the LLM backbone frozen. We introduce a VKD scheme, which incorporates intermediate visual representation from a teacher model as a supervision signal. Interestingly, VKD not only facilitates forgetting of target visual knowledge but also aids in preserving visual knowledge of non-target entities. Meanwhile, we combine neuron pruning and FIM masking in a unified framework. Moreover, we evaluate the robustness of our MLLM unlearning approach against relearning attacks, demonstrating that the forgotten visual knowledge cannot be easily recovered.